OpenStudy (anonymous):

A swimming pool has to be drained for maintenance. The pool is shaped like a cylinder with a diameter of 10 m and a depth of 1.7 m . If the water is pumped out of the pool at the rate of 14 m3 per hour, how many hours does it take to empty the pool? Use the value 3.14 for π, and round your answer to the nearest hour.

OpenStudy (anonymous):

whats the reward

OpenStudy (anonymous):

idk whatever you want?

OpenStudy (anonymous):

i just need help ASAP

OpenStudy (anonymous):

porn?

OpenStudy (anonymous):

yeahhh NO! I'm not that desperate!

OpenStudy (anonymous):

i know I was joking

OpenStudy (anonymous):

okay....

OpenStudy (anonymous):

its 13.8

OpenStudy (anonymous):

no it's not!

OpenStudy (anonymous):

what can it be then cus thats what I came up with

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i got 6 hours

OpenStudy (anonymous):

no its wrong! Thank you tho

OpenStudy (anonymous):

whats wrong

OpenStudy (anonymous):

the answer lol

OpenStudy (anonymous):

Round each number to the nearest hundredth. What is the best estimate for the answer to 434.657 + 102.494? A. 536.05 B. 537.04 C. 537.15 D. 537.16

OpenStudy (anonymous):

@HerbvanH need ur help

OpenStudy (anonymous):

idk then bc i check 3x and i got 5.6

OpenStudy (anonymous):

its cool thank you!

OpenStudy (anonymous):

Round each number to the nearest hundredth. What is the best estimate for the answer to 434.657 + 102.494? A. 536.05 B. 537.04 C. 537.15 D. 537.16

OpenStudy (anonymous):



OpenStudy (anonymous):

hey beaut check ur pm

OpenStudy (anonymous):

so, the contents = \[\pi \times diameter \times h\]

OpenStudy (anonymous):

53,35 / 14 = 3,81285714..... = 4